A bar magnet is oriented with its magnetic dipole moment at 30.7° to the direction of a 0.0107-T magnetic field, causing a torque of 0.00487 N·m to act on it. Find the magnitude of the magnet\'s magnetic dipole moment.
Answer #1

(torque) = MB sin@. we can get the required p.
M- the magnetic dipole moment. B-magnetic field induction in tesla and @-the angle between the field and the dipole moment
So M= tau/B sin@ = 0.00487/0.0107*sin30.7 = 0.8914 Am^2